# A Wireless AI-Generated Content (AIGC) Provisioning Framework Empowered by Semantic Communication

### v2

@@ Abstract
Generative AI applications have been recently catering to a vast user base by creating diverse and high-quality AI-generated content (AIGC). With the proliferation of mobile devices and rapid growth of mobile traffic, providing ubiquitous access to high-quality AIGC services via wireless communication networks is becoming the future direction. However, it is challenging to provide qualified AIGC services in wireless networks with unstable channels, limited bandwidth resources, and unevenly distributed computational resources. To tackle these challenges, we propose a semantic communication (SemCom)-empowered AIGC (SemAIGC) generation and transmission framework, where only semantic information of the content rather than all the binary bits should be generated and transmitted by using SemCom. Specifically, SemAIGC integrates diffusion models within the semantic encoder and decoder to design a workload-adjustable transceiver thereby allowing adjustment of computational resource utilization in edge and local. In addition, a resource-aware workload trade-off (ROOT) scheme is devised to intelligently make workload adaptation decisions for the transceiver, thus efficiently generating, transmitting, and fine-tuning content as per dynamic wireless channel conditions and service requirements. Simulations verify the superiority of our proposed SemAIGC framework in terms of latency and content quality compared to conventional approaches.
@@ I Introduction
Driven by recent advancements in deep learning and computing hardware, generative AI has demonstrated remarkable strides in analyzing various forms of media and creating AI-generated content (AIGC) [1]. For the delivery of high-quality content, well-crafted AIGC models have evolved to maintain millions of parameters within their neural networks. This feat can necessitate a large-scale server devouring thousands of hours of graphics processing unit (GPU) time in a single model training session [2]. Limited by the computing-intensive nature of AIGC, existing AIGC applications, including notable products like ChatGPT, DELL-R-2, and ERNIE Bot, predominantly find their home on cloud servers equipped with ample computational resources [3].
I-A Background: AIGC in Wireless
@@
Due to the rapid development of mobile technology and the increasing adoption of smartphones, tablets, and other mobile devices, the percentage of global Internet traffic on wireless devices has surged over the past decade. In 2022, it was estimated that there were over 15 billion wireless devices worldwide, which contributed 55% of global Internet traffic [4]. In this background, many AI companies are actively strategizing and positioning themselves in the realm of wireless device-oriented AIGC applications. However, wireless AIGC services demand strict requirements on content quality and latency, which pose significant challenges under dynamic wireless channels. Moreover, these services often grapple with limited communication resources (like bandwidth, and high-quality channels) and unevenly distributed computational resources in wireless networks. Therefore, more efficient and accessible AIGC generation and transmission frameworks are required for wireless AIGC service provisioning.
@@
Recently, a few studies about wireless AIGC generation and transmission have been conducted to improve the utilization of unevenly distributed computational resources. The authors in [5] propose a wireless AIGC network architecture where cloud servers primarily handle computationally intensive AIGC model pre-training. Additionally, it offloads content generation tasks from cloud servers to edge transmitters. In [6], a dynamic AIGC service provider selection scheme is proposed in an intelligent wireless AIGC network to enable users to connect to the provider with suitable edge servers and enough computational resources. Considering the limited bandwidth resources and dynamic channel quality, the authors in [7] develop a pricing-based incentive mechanism for AIGC generation and transmission framework to maximize users’ utility in mobile edge networks. These existing works didn’t address the insufficient communication resources in wireless networks when coping with the potentially overwhelming number of service requests.
@@
In parallel, some research works aim to deploy AIGC models in local devices to address the communication resource shortage in the wireless network. Lightweight AIGC models have been tested and certified in works [8, 9], thus it is becoming possible to deploy AIGC models on local mobile devices with limited computational resources. Under this premise, a novel collaborative distributed AIGC framework is conceptualized in [10], which allows smartphones to process the diffusion model locally in an edge server-empowered wireless AIGC network. In this way, less content is transmitted from edge to local, thus the pressure of communication resources is relieved. However, considering the computational resource limitation of mobile devices, the generated content quality and content-generating latency are difficult to ensure.
I-B Motivation
@@
To explore an AIGC service provisioning framework in wireless networks, it is required to meet diverse service requirements under limited communication resources, unevenly distributed computational resources, and dynamic channel quality. In this case, semantic communication (SemCom) is a potential technique to empower wireless AIGC provisioning. SemCom, aims to effectively transmit necessary semantic information of source message, as opposed to the accurate reception of every single binary bit without considering its meaning [11, 12, 13, 14]. Empowered by SemCom, the generated semantic information from the AIGC model can be transmitted instead of the entire bits of content, and the pressure of communication resources, especially bandwidth, can be relieved with a slim data size [15]. Meanwhile, the receiver accurately interprets semantic information from transmitted bits and reduces semantic distortion caused by channel noise, thus improving transmission reliability at a semantic level and allowing for suitable AIGC services over harsh wireless channels [16]. With these superiorities, combining SemCom with generative AI models is promising to provide AIGC service over wireless networks with limited communication resources and unevenly distributed computational resources.
@@
However, resources in wireless networks, such as bandwidth, GPU/CPU cores, and computational power, are often with dynamic availability. Therefore, it is inefficient to directly integrate SemCom and AIGC with a fixed encoder and decoder structure, where the AIGC model is merely deployed within the encoder side or decoder side. This fixed structure makes the encoder/decoder unable to adjust the computing workloads according to the resource availability. Furthermore, in the fixed structure of the encoder and decoder, the semantic density is unadjustable, causing a high semantic noise ratio under bad channel conditions, thereby posing additional denoising difficulty in decoding. Meanwhile, users may have varying service requirements, such as latency and content quality, when accessing various AIGC services. Since the semantic computing workloads and semantic density are unadjustable in the transceiver, meeting these requirements becomes another difficulty.
@@
In response to the unsolved difficulties, deploying diffusion models in both edge and local, and employing machine learning (ML) in the joint AIGC generation and transmission is a potential solution. The diffusion models, as popular generative AI models, are utilized to denoise irrelevant information from pure Gaussian noise and finally retain valid information of content like text [17], audio [18], or image [19]. By deploying diffusion models into the encoder and decoder, it is promising to develop a workload-adjustable transceiver, in which the semantic information can be generated within both edge and local. Simultaneously, ML is capable of intelligently deciding the denoising steps of the encoder and decoder as per resource availability and channel quality, i.e., optimizing the workload adaptation, thus meeting the content generation latency requirement [20]. Furthermore, semantic density can also be optimized to reduce the semantic noise ratio, thus efficiently ensuring that the content quality is consistent and satisfies the requirements under varying channel qualities.
I-C Contributions and Organization
@@
  • We theoretically prove that the modified diffusion model deployed at the receiver to can denoise the semantic noise caused by the wireless channel.
@@
This paper is organized as follows. In Section II, we present the SemAIGC framework, followed by the design of edge transmitter and local receiver in Section III. The ROOT scheme design for the workload-adjustable transceiver is elaborated in Section IV. Then, we evaluate the performance gain of the SemAIGC framework and ROOT scheme in Section V. Section VI concludes the paper.
@@
II SemCom-Empowered AIGC Generation and Transmission Framework
@@
II-A Cloud Pre-Training
In SemAIGC, we utilize a cloud server to conduct the text-to-image (T2I) model training processes. Generally, a high-quality T2I model requires millions of parameters in a total size of 1-10 GB, stringent requirements are exposed for specialized hardware like GPU, CPU, and capable devices during its training. The cloud server as a centralized infrastructure with ample computing power, storage, database, etc. can be well-matched to the demands. As shown in Fig. 1, the cloud server retains the copy of all the modules within both the encoder and decoder. After feeding these modules with a substantial corpus of text-image pairs, the edge transmitter and local receiver can download the pre-trained encoder and decoder from the cloud server, respectively. Additionally, the encoder and decoder can be periodically updated within the cloud server, while the edge transmitter and local receiver do not require further training before using them to offer AIGC access services. Moreover, the neural network of the ROOT scheme is also trained remotely and later deployed in the edge transmitter, then fine-tuned according to the local wireless communication environment.
II-B Edge Encoding
As shown in Fig. 1, the edge transmitter is composed of a semantic extraction module, a semantic generation module, and a channel encoder. The semantic extraction module comprises two independent networks, responsible for extracting semantic information from text and image, respectively. As the core of the edge transmitter, the semantic generation module uses a diffusion model to predict and gradually denoise data starting from pure noise according to the provided textual description. Compared with the pre-training stage, fewer computational resources are required in the AIGC generation and transmission stage. Therefore, the transmitter with edge computing servers is sufficient to run these modules with acceptable computing latency. Apart from the semantic encoder, the channel encoder in Fig. 1 is deployed at the edge transmitter to convert semantic information into binary bits for transmission.
II-C Local Decoding
With the improvement of computing power in mobile devices, plenty of user devices are capable of executing information fine-tuning and recovery processes. In SemAIGC, the T2I decoder of a local receiver in Fig. 1 is composed of a semantic fine-tuning module, an autoencoder decoder module, and a channel decoder. Specifically, the channel decoder converts the binary bits into semantic information. Given the limited computing power of local devices, a lightweight diffusion model is deployed into the semantic fine-tuning module. After a rapid fine-tuning process, semantic noise can be eliminated, and then the autoencoder decoder module restructures semantic information into high-resolution images with text guidance.
@@ II-C Local Decoding
III Workload-Adjustable Transceiver Design
@@
III-A Encoder Design at Edge Transmitter
@@
1) Semantic extraction module: Let ${s}_{\scriptscriptstyle t\!e\!x\!t}$ and ${s}$ denote the input text and image to the semantic extraction module, respectively. The semantic information extracted from text and that from image are denoted as $z_{\scriptscriptstyle t\!e\!x\!t}$ and $z$, respectively. As the guidance of image generation, the semantic information of text $z_{\scriptscriptstyle t\!e\!x\!t}$ is extracted by a contrastive language-image pre-training (CLIP) encoder [21] as a sequence. Meanwhile, by using an attention-based variational autoencoder (VAE) model [22], the input image is encoded into Gaussian-like distribution and outputs multiple sets of means and standard deviations. Then, we use these values to generate latent space image $z$, i.e., image semantic information. Note that the semantic information of real image is only used during the model training stage and is not directly involved in the T2I generation. The two modalities’ semantic information is extracted as follows:
$z_{\scriptscriptstyle t\!e\!x\!t}=\mathcal{E}\left({s}_{\scriptscriptstyle t\!e\!x\!t};{\varphi}_{\scriptscriptstyle t\!e\!x\!t}\right),{z}=\mathcal{E}\left({s};\varphi\right).$ (1)
Here, $\mathcal{E}\left(\cdot;{\varphi}_{\scriptscriptstyle t\!e\!x\!t}\right)$ and $\mathcal{E}\left(\cdot;\varphi\right)$ are the text encoder and image encoder with learnable parameters ${\varphi}_{\scriptscriptstyle t\!e\!x\!t}$ and ${\varphi}$, respectively.
@@
The semantic generation modules in Fig. 2 is based on the denoising diffusion implicit models (DDIMs), consisting of the diffusion process and the reverse diffusion process [23].
@@
As shown in Fig. 3, the diffusion process is a fixed (or predefined) forward diffusion process, which is denoted as $q(z_{t}|z_{t-1})$. In the diffusion process, a scheduler gradually adds Gaussian noise at each time step $t\in[0,T]$, until the initial semantic information of an image $z_{0}$ becomes pure noise $z_{\scriptscriptstyle T}$. Basically, each new image at time step $t$ is drawn from a conditional Gaussian distribution as
$z_{t}=\sqrt{1-\beta_{t}}z_{t-1}+\sqrt{\beta_{t}}\epsilon,$ (2)
where $10<\beta_{1}<\beta_{2}<\cdots<\beta_{\scriptscriptstyle T}<$ 1 is a known variance schedule with time-related constants, and $\epsilon\sim{\cal N}(0,I)$ is the added Gaussian noise with an identity matrix $I$.
@@
Meanwhile, the reverse diffusion process in Fig. 3 is a denoising process which is represented as $p(z_{t-1}|z_{t})$. In this process, a neural network, i.e., UNet [24], is trained to learn the conditional probability distribution $p_{\theta}(z_{t-1}|z_{t},z_{\scriptscriptstyle t\!e\!x\!t})$ and gradually denoise the semantic information according to text semantic starting from pure noise until it ends up with the semantic information of an actual image. Here, $\theta$ is the learnable parameters of the UNet, text guidance is implemented by concatenating the text embedding to the key-value pairs of each self-attention layer in the UNet [19]. Accordingly, the reverse diffusion process is
@@
$z_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(z_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\dot{\alpha_{t}}}}\epsilon_{\theta}\left(z_{t},t,z_{\scriptscriptstyle t\!e\!x\!t}\right)\right)+\bar{\sigma}_{t}\tilde{\epsilon},$ (3)
@@
where $\epsilon_{\theta}$ is the UNet predicted noise, $\alpha_{t}=1-\beta_{t}$ is a known constant of step $t$, $\dot{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{i}$ is a cumulative product of $\alpha_{t}$, $\bar{\sigma}_{t}$ is a time dependent constant of step $t$, and $\tilde{\epsilon}\sim{\cal N}(0,I)$ is the random normal Gaussian noise.
@@
Note that the semantic generation module conducts different processes during the pre-training stage and content-generating stage, as shown in Fig. 3. In the pre-training stage, the semantic generation module executes both processes to learn the noise distribution, while only the reverse diffusion process is needed in content-generating stage. In the content-generating stage, with pre-trained UNet, the target image is denoised from randomly generated pure noisy data $z_{\scriptscriptstyle T}$. After the scheduled denoising steps $T$, the image semantic information is output according to text guidance, which is expressed as
$z_{\theta}=\mathcal{F}_{1}(z_{\scriptscriptstyle T},T,z_{\scriptscriptstyle t\!e\!x\!t};{\theta}),$ (4)
@@
where $\mathcal{F}_{1}(\cdot;\theta)$ is the UNet with learnable parameter $\theta$. The model training of diffusion models will be elaborated in Section III-C.
@@
3) Channel encoder module: As very few bits are required for text semantic information transmission, let us only focus on the transmission of image semantic information [25]. The transmitted signal is encoded as ${y}=\mathcal{C}({z_{\theta}})$ via the channel encoder.
@@
${y}^{\prime}=h{y}+n,$ (5)
where $h$ represents the channel gain and $n$ is the noise. We consider the additive white Gaussian noise (AWGN) channel in this paper. [^1] The used diffusion model is Gaussian noise-based hot diffusion, which is designed to denoise Gaussian distributed noise. For filtering other types of noise, like Rayleigh noise, a cold diffusion-based module could be necessary [26].
III-B Decoder Design at Local Receiver

### v3

@@ Abstract
With the significant advances in AI-generated content (AIGC) and the proliferation of mobile devices, providing high-quality AIGC services via wireless networks is becoming the future direction. However, the primary challenges of AIGC services provisioning in wireless networks lie in unstable channels, limited bandwidth resources, and unevenly distributed computational resources. To this end, this paper proposes a semantic communication (SemCom)-empowered AIGC (SemAIGC) generation and transmission framework, where only semantic information of the content rather than all the binary bits should be generated and transmitted by using SemCom. Specifically, SemAIGC integrates diffusion models within the semantic encoder and decoder to design a workload-adjustable transceiver thereby allowing adjustment of computational resource utilization in edge and local. In addition, a resource-aware workload trade-off (ROOT) scheme is devised to intelligently make workload adaptation decisions for the transceiver, thus efficiently generating, transmitting, and fine-tuning content as per dynamic wireless channel conditions and service requirements. Simulations verify the superiority of our proposed SemAIGC framework in terms of latency and content quality compared to conventional approaches.
@@ 1 Introduction
Generative AI has demonstrated remarkable strides in analyzing various forms of media and creating AI-generated content (AIGC) [1]. For the delivery of high-quality content, well-crafted AIGC models have evolved to maintain millions of parameters within their neural networks. This feat can necessitate a large-scale server devouring thousands of hours of graphics processing unit (GPU) time in a single model training session [2]. Limited by the computing-intensive nature of AIGC, existing AIGC applications, including notable products like ChatGPT, DELL-R-2, and ERNIE Bot, predominantly find their home on cloud servers equipped with ample computational resources [3].
1.1 Background: AIGC in Wireless
@@
Driven by the ubiquitous wireless connectivity and surge global traffic of terminal devices [4], many AI companies are actively strategizing and positioning themselves in the realm of wireless device-oriented AIGC applications. However, wireless AIGC services demand strict requirements on content quality and latency, which pose significant challenges under dynamic wireless channels. Moreover, these services often grapple with limited communication resources (like bandwidth, and high-quality channels) and unevenly distributed computational resources in wireless networks. Therefore, it calls urgently for efficient and accessible AIGC generation and transmission frameworks underpinned by wireless networks.
@@
Recently, a few studies about wireless AIGC generation and transmission have been conducted to improve the utilization of unevenly distributed computational resources, as listed in Table I. The authors in [5] propose a wireless AIGC network architecture where cloud servers primarily handle computationally intensive AIGC model pre-training. Additionally, it offloads content generation tasks from cloud servers to edge transmitters. In [6], a dynamic AIGC service provider selection scheme is proposed in an intelligent wireless AIGC network to enable users to connect to the provider with suitable edge servers and enough computational resources. Considering the limited bandwidth resources and dynamic channel quality, the authors in [7] develop a pricing-based incentive mechanism for AIGC generation and transmission framework to maximize users’ utility in mobile edge networks. These existing works didn’t address the insufficient communication resources in wireless networks when coping with the potentially overwhelming number of service requests.
@@
In parallel, some research works aim to deploy AIGC models in local devices to address the communication resource shortage in the wireless network, as listed in Table I. Lightweight AIGC models have been proposed and tested in works [8, 9]. Moreover, some experiments validate the capability of mobile devices for AIGC model deployment, for example, an image can be generated within 2 seconds on a typical smartphone [10]. Under this premise, a novel collaborative distributed AIGC framework is conceptualized in [11], which allows smartphones to process the diffusion model locally in an edge server-empowered wireless AIGC network. In this way, less content is transmitted from edge to local, thus the pressure of communication resources is relieved. However, considering the computational resource limitation of mobile devices, the generated content quality and content-generating latency are difficult to ensure.
TABLE I: Comparison of different AIGC frameworks
AIGC Frameworks Features Related Literatures
Conventional Cloud AIGC These models run on remote servers, generally with high content quality and low computation latency, but lead to high transmission latency, and high core network access latency [1],[2],[3]
Edge and Local AIGC Deploying AIGC models in edge servers, and local devices, flexibly choosing suitable edge servers or local devices to generate contents as per communication resource and computational ability. [5],[6],[7],[8],[9],[10],[11]
SemCom-empowered AIGC Integrating SemCom into the AIGC model and deploying modularized components at both the transmitter and receiver, thus enabling the transmission of semantic information between them, and allowing edge servers and local devices to collaborate in content generation using distributed computational resources. [12]
1.2 Motivation
@@
To explore an AIGC service provisioning framework in wireless networks, it is required to meet diverse service requirements under limited communication resources, unevenly distributed computational resources, and dynamic channel quality. In this case, semantic communication (SemCom) is a potential technique to empower wireless AIGC provisioning. SemCom aims to effectively transmit necessary semantic information of source message, as opposed to the accurate reception of every single binary bit without considering its meaning [13, 14, 15, 16, 17]. In this way, the pressure of communication resources, especially bandwidth, can be relieved with a slim data size [18]. Meanwhile, the receiver with semantic decoder accurately interprets semantic information from distorted bits, thus improving transmission reliability at a semantic level and allowing for suitable AIGC services over harsh wireless channels [19]. Furthermore, SemCom models and generative AI models usually exploit similar encoder and decoder structures and could involve overlapping neural network layers, which makes it natural to integrate these models without introducing additional system complexity [20].
@@
However, resources in wireless networks, such as bandwidth, GPU/CPU cores, and computational power, are often with dynamic availability. Therefore, it is inefficient to directly integrate SemCom and AIGC with a fixed encoder and decoder structure, where the AIGC model is merely deployed within the encoder side or decoder side. This fixed structure makes the encoder/decoder unable to adjust the computing workloads according to the resource availability. Furthermore, in the fixed structure of the encoder and decoder, the semantic density is unadjustable, causing a high semantic noise ratio under bad channel conditions, thereby posing additional denoising difficulty in decoding [21]. Meanwhile, users may have varying service requirements, such as latency and content quality, when accessing various AIGC services. Since the semantic computing workloads and semantic density are unadjustable in the transceiver, meeting these requirements becomes another difficulty.
@@
In response to the unsolved difficulties, deploying diffusion models in both edge and local for the joint AIGC generation and transmission is a potential solution. The diffusion models, as popular generative AI models, are utilized to denoise irrelevant information from pure Gaussian noise and finally retain valid information of content like text [22], audio [23], or image [24]. By deploying diffusion models into the encoder and decoder, it is promising to develop a workload-adjustable transceiver, in which the semantic information can be generated within both edge and local. Additionally, machine learning (ML) is capable of intelligently deciding the denoising steps of the encoder and decoder as per resource availability and channel quality, i.e., optimizing the workload adaptation, thus meeting the content generation latency requirement. Furthermore, semantic density can also be optimized to reduce the semantic noise ratio, thus efficiently ensuring that the content quality is consistent and satisfies the requirements under varying channel conditions.
1.3 Contributions and Organization
@@
  • We theoretically prove that the modified diffusion model deployed at the receiver to denoise the semantic noise caused by the wireless channel.
@@
This paper is organized as follows. In Section 2, we present the SemAIGC framework, followed by the design of edge transmitter and local receiver in Section 3. The ROOT scheme design for the workload-adjustable transceiver is elaborated in Section 4. Then, we evaluate the performance gain of the SemAIGC framework and ROOT scheme in Section 5. Section 6 concludes the paper.
@@
2 SemCom-Empowered AIGC Generation and Transmission Framework
@@
2.1 Cloud Pre-Training
In SemAIGC, we utilize a cloud server to conduct the text-to-image (T2I) model training processes. Generally, a high-quality T2I model requires millions of parameters in a total size of 1-10 GB, stringent requirements are exposed for specialized hardware like GPU, CPU, and capable devices during its training. The cloud server as a centralized infrastructure with ample computing power, storage, database, etc. can be well-matched to the demands. As shown in Fig. 1, the cloud server retains the copy of all the modules within both the encoder and decoder. After feeding these modules with a substantial corpus of text-image pairs, the edge transmitter and local receiver can download the pre-trained encoder and decoder from the cloud server, respectively. Since the large models within the encoder-decoder generally perform stable after the initial offline pre-training, communication overhead could be very minor. In addition, even if periodical updates of the encoder and decoder are required, these models can be modularized as shown in Fig. 1, and each update might merely cover partial modules and lead to less data transmission. Moreover, the neural network of the ROOT scheme can also be trained remotely and later deployed in the edge transmitter, then fine-tuned according to the local wireless network environment.
2.2 Edge Encoding
As shown in Fig. 1, the edge transmitter is composed of a semantic extraction module, a semantic generation module, and a channel encoder. The semantic extraction module comprises two independent networks, responsible for extracting semantic information from text and image, respectively. As the core of the edge transmitter, the semantic generation module uses a diffusion model to predict and gradually denoise data starting from pure Gaussian noise according to the provided textual description. Compared with the pre-training stage, fewer computational resources are required in the AIGC generation and transmission stage. Therefore, the transmitter with edge computing servers is sufficient to run these modules with acceptable computing latency. Apart from the semantic encoder, the channel encoder in Fig. 1 is deployed at the edge transmitter to convert semantic information into binary bits for transmission.
2.3 Local Decoding
With the improvement of computing power in mobile devices, plenty of user devices are capable of executing information fine-tuning and recovery processes. In SemAIGC, the T2I decoder of a local receiver in Fig. 1 is composed of a semantic fine-tuning module, an autoencoder decoder module, and a channel decoder. Specifically, the channel decoder converts the binary bits into semantic information. Given the limited computing power of local devices, a lightweight diffusion model is deployed into the semantic fine-tuning module, which has similar functions as the aforementioned semantic generation module. After a rapid fine-tuning process, semantic noise can be eliminated, and then the autoencoder decoder module restructures semantic information into high-resolution images with text guidance.
@@ 2.3 Local Decoding
3 Workload-Adjustable Transceiver Design
@@
3.1 Encoder Design at Edge Transmitter
@@
1) Semantic extraction module: Let $\bm{s}_{\scriptscriptstyle t\!e\!x\!t}\in\mathbb{R}^{D}$ and $\bm{s}\in\mathbb{R}^{D_{H}\times D_{W}\times D_{C}}$ denote the input text and image to the semantic extraction module, respectively. Here, $D$ is the number of tokens of input text $\mathbf{s}_{\scriptscriptstyle t\!e\!x\!t}$, and $D_{H}\times D_{W}\times D_{C}$ are the dimensions for a color image with $D_{C}$ the number of channels (e.g., 3 for the RGB image). Then, the semantic information extracted from the text and that from the image are denoted as $\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t}\in\mathbb{R}^{d}$ and $\mathbf{z}\in\mathbb{R}^{d_{h}\times d_{w}\times d_{c}}$, respectively. Here, text semantic information is a vector of dimension $d$, and $d_{h}\times d_{w}\times d_{c}$ are the dimensions for image semantic information, typically $d_{h}\ll D_{H},d_{w}\ll D_{W}$. As the guidance of image generation, the semantic information of text $\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t}$ is extracted by a contrastive language-image pre-training (CLIP) encoder [25]. Meanwhile, by using an attention-based variational autoencoder (VAE) model [26], the input image is encoded into the Gaussian-like distribution and outputs multiple sets of means and standard deviations. Then, we use these values to generate latent space image $\mathbf{z}$, i.e., image semantic information. Note that the semantic information of real images is only used during the model training stage and is not directly involved in the T2I generation. The two modalities’ semantic information is extracted as
$\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t}=\mathcal{E}\left(\mathbf{s}_{\scriptscriptstyle t\!e\!x\!t};\bm{\varphi}_{\scriptscriptstyle t\!e\!x\!t}\right),\mathbf{z}=\mathcal{E}\left(\mathbf{s};\bm{\varphi}\right).$ (1)
Here, $\mathcal{E}\left(\cdot;\bm{\varphi}_{\scriptscriptstyle t\!e\!x\!t}\right)$ and $\mathcal{E}\left(\cdot;\bm{\varphi}\right)$ are the text encoder and image encoder with learnable parameters $\bm{\varphi}_{\scriptscriptstyle t\!e\!x\!t}$ and $\bm{\varphi}$, respectively.
@@
The semantic generation modules in Fig. 2 is based on the denoising diffusion probabilistic models (DDPMs), consisting of the diffusion process and the reverse diffusion process [27]. As shown in Fig. 3, the diffusion process is a fixed (or predefined) forward diffusion process, which is denoted as $q(\mathbf{z}_{t}|\mathbf{z}_{t-1})$. In the diffusion process, a scheduler gradually adds Gaussian noise at each time step $t\in[0,T]$, until the initial semantic information of an image $\mathbf{z}_{0}$ becomes pure noise $\mathbf{z}_{\scriptscriptstyle T}$, where $\mathbf{z}_{t}\in\mathbb{R}^{d_{h}\times d_{w}\times d_{c}}$. Therefore, the diffusion process is
$\mathbf{z}_{t}=\sqrt{1-\beta_{t}}\mathbf{z}_{t-1}+\sqrt{\beta_{t}}\bm{\epsilon},$ (2)
@@
where $10<\beta_{1}<\beta_{2}<\cdots<\beta_{\scriptscriptstyle T}<$ 1 is a known variance schedule with time-related constants, and $\bm{\epsilon}\sim{\cal N}(0,\bm{I}),\bm{\epsilon}\in\mathbb{R}^{d_{h}\times d_{w}\times d_{c}}$ is the added Gaussian noise with an identity matrix $\bm{I}$. Meanwhile, the reverse diffusion process in Fig. 3 is a denoising process which is represented as $p(\mathbf{z}_{t-1}|\mathbf{z}_{t})$. In this process, a neural network, i.e., UNet [28], is trained to learn the conditional probability distribution $p_{\bm{\theta}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t})$ and gradually denoise the semantic information according to text semantic starting from pure noise until it ends up with the semantic information of an actual image. Here, $\bm{\theta}$ is the learnable parameters of the UNet, text guidance is implemented by concatenating the text embedding to the key-value pairs of each self-attention layer in the UNet [24]. Accordingly, the reverse diffusion process is
@@
$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\dot{\alpha_{t}}}}\bm{\epsilon}_{\bm{\theta}}\left(\mathbf{z}_{t},t,\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t}\right)\right)+\bar{\sigma}_{t}\tilde{\bm{\epsilon}},$ (3)
@@
where $\bm{\epsilon}_{\bm{\theta}}$ is the UNet predicted noise, $\alpha_{t}=1-\beta_{t}$ is a known constant of step $t$, $\dot{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{i}$ is a cumulative product of $\alpha_{t}$, $\bar{\sigma}_{t}$ is a time dependent constant of step $t$, and $\tilde{\bm{\epsilon}}\sim{\cal N}(0,\bm{I})$ is the random normal Gaussian noise.
@@
Note that the semantic generation module conducts different processes during the pre-training stage and content-generating stage, as shown in Fig. 3. In the pre-training stage, the semantic generation module executes both processes to learn the noise distribution, while only the reverse diffusion process is needed in content-generating stage. In the content-generating stage, with pre-trained UNet, the target image is denoised from randomly generated pure noisy data $\mathbf{z}_{\scriptscriptstyle T}$. After the scheduled denoising steps $T$, the image semantic information is output according to text guidance, which is expressed as
$\mathbf{z}_{\bm{\theta}}=\mathcal{F}_{1}(\mathbf{z}_{\scriptscriptstyle T},T,\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t};{\bm{\theta}}),$ (4)
@@
where $\mathcal{F}_{1}(\cdot;\bm{\theta})$ is the UNet with learnable parameter $\bm{\theta}$. The model training of diffusion models will be elaborated in Section 3.3.
@@
3) Channel encoder module: As very few bits are required for text semantic information transmission, let us only focus on the transmission of image semantic information [29]. The transmitted signal is encoded as $\mathbf{Y}=\mathcal{C}({\mathbf{z}_{\bm{\theta}}})$ via the channel encoder, where $\mathbf{Y}\in\mathbb{C}^{M_{c}\times k}$ is the analog signal, $k$ denotes the number of channels used to transmit the signal.
@@
$\mathbf{Y}^{\prime}=\mathbf{H}\mathbf{Y}+\mathbf{N},$ (5)
where $\mathbf{H}\in\mathbb{C}^{M_{c}\times M_{c}}$ represents the channel gain matrix and $\mathbf{N}$ is the noise. We consider the additive white Gaussian noise (AWGN) channel in this paper. [^1] The used diffusion model is Gaussian noise-based hot diffusion, which is designed to denoise Gaussian distributed noise. For filtering other types of noise, like Rayleigh noise, it is necessary to specifically pre-train a cold diffusion-based module [30].
3.2 Decoder Design at Local Receiver